OpenStudy (anonymous):

63(p^2 -1)=32p by factoring

OpenStudy (venomblast):

difference of perfect square (p-1)(p+1)

OpenStudy (venomblast):

x=-7/9 or x=9/7

OpenStudy (anonymous):

Thanks that's what i thought it was but i wanted to be sure.

OpenStudy (venomblast):

hehe no problem